# Enhancing Human-Robot Collaborative Assembly in Manufacturing Systems Using Large Language Models

### v1

@@ IV-E Case Study Discussion and Limitations
The case study evaluated the system’s ability to facilitate collaboration between humans and robots and showcases how the integration of LLMs can lead to efficient and flexible manufacturing processes. The results demonstrate that as the instructions become less specific, the robot’s performance decreases significantly, indicating that the robot agent needs well-defined human commands to operate effectively. These results highlight the limitations of the proposed framework and suggest an area for improvement. One such improvement involves introducing feedback mechanisms that allow the system to ask for clarification when instructions are unclear, refining the interaction between humans and robots. Notably, the robot’s success was higher with commands that included keywords like ’fixed’ or ’addressed,’ which means that the robot understands its initialization. The initialization protocol informs the robot that the human’s role is to correct errors to resume the assembly task. Based on the case study, we have found that the initialization process is crucial as it ensures that the robot can understand the roles within the system.
There were also limitations in the case study and evaluation of the proposed framework. Specifically, we only evaluated a limited range of commands from the operator. These commands were specifically related to predefined assembly scenarios and did not include other interactions, such as human interruptions and task-irrelevant questions from the human. The case study also did not study the variability in the operator’s knowledge, as the framework assumed operators would provide relevant instructions to the assembly task. Furthermore, the operator was not allowed to change the defined tasks, e.g., variations in positioning, or task order. Future work will focus on analyzing the developed framework in a variety of manufacturing scenarios. We will look to test the proposed system with operators who do not have prior knowledge of the assembly process. We will also test the proposed framework for other material handling and manufacturing tasks.
@@ V Conclusion and Future Work
Advancements in LLM are applied to human-robot collaborative assembly for executing robot actions and collaborating with humans based on environmental data. By incorporating LLM, robots can better understand and interact with human operators to resolve errors and improve execution based on environmental feedback. Drawing on insights from these studies, we incorporated LLM into our human-robot collaborative assembly framework, enabling dynamic responses to task variations.
This research addresses key challenges in human-robot collaborative assembly, including developing communication systems that require minimal robotics training (C1), enhancing adaptability and flexibility to manage real-time changes and errors in human-robot collaborative assembly within the manufacturing environment (C2), and integrating advanced technologies into human-robot collaborative assembly process with a human-centric design to improve usability and communication (C3). To validate the effectiveness of our proposed framework, we implemented a practical setup involving the cable shark device assembly process. This setup showcased the framework’s capability to facilitate intuitive human-robot communication through voice commands with language variations. We dynamically adapt to task variations and errors by integrating LLM, sensors, and task control mechanisms, demonstrating its ability to maintain productivity and ensure a continuous workflow.
Future work includes extending this LLM-based human-robot collaborative assembly framework by focusing on increasing its adaptability by feeding LLM with a wide range of data, such as various robotic tasks and sensor data. The aim is to increase the robot’s flexibility in performing tasks, enhance safety, and enable it to address unexpected errors and task variations. Additionally, we also plan to incorporate multiple modalities, such as haptic and gestures to improve human-robot interaction within the manufacturing environment. Our framework plans to enhance this multimodal strategy by integrating LLM to improve communication efficiency, task adaptability, and intuitive interaction between humans and robots within the manufacturing environment.

### v2

@@ IV-E Case Study Discussion and Limitations
The case study evaluated the system’s ability to facilitate collaboration between humans and robots and showcases how the integration of LLMs can lead to efficient and flexible manufacturing processes. The results show that as instructions become less specific, robot performance significantly decreases, indicating the need for well-defined commands. For example, the ambiguous command "Correction is made. Resume the operations." failed due to lack of context and explicit task reference. These results highlight the limitations of the proposed framework and suggest an area for improvement. One such improvement involves introducing feedback mechanisms that allow the system to ask for clarification when instructions are unclear, refining the interaction between humans and robots. Another approach is to fine-tune the LLM [29] or integrate a knowledge base [30] with specific manufacturing processes to improve its ability to handle instructions and provide accurate responses. Commands containing keywords like ’fixed’ or ’addressed’ had higher success rates, emphasizing the importance of initialization protocol. The protocol informs the robot that the human’s role is to correct errors to resume the assembly task. The case study found that a clear initialization process is crucial for the robot to understand its role and functions within the system.
There were also limitations in the case study and evaluation of the proposed framework. Specifically, we only evaluated a limited range of commands from the operator, rather than assessing the overall system capabilities. These commands were related to predefined assembly scenarios and did not include other interactions, such as human interruptions and task-irrelevant questions from the human. While LLM-generated error messages were understandable for knowledgeable operators, there is a risk of low-quality messages for less experienced operators or in varied contexts. The case study also did not study the variability in the operator’s knowledge, as the framework assumed operators would provide relevant instructions to the assembly task. Furthermore, the operator was not allowed to change the defined tasks, e.g., variations in positioning or task order.
Future work will analyze the developed framework in various manufacturing scenarios. We will involve users without assembly knowledge, allowing free interaction to complete tasks. Additionally, we will compare the system to other methods by evaluating adaptability, error handling, and performance. The framework will also be assessed using different initialization protocols and tested on various manufacturing assembly tasks.
@@ V Conclusion and Future Work
Advancements in LLM are applied to human-robot collaborative assembly to execute actions and collaborate based on environmental data. Incorporating LLM allows robots to better understand human operators, resolve errors, and improve execution with environmental feedback. Based on these insights, we integrated LLM into our assembly framework for dynamic responses to task variations.
This research addresses key challenges in human-robot collaborative assembly, including developing communication systems that require minimal robotics training (C1), enhancing adaptability to manage changes and errors (C2), and integrating advanced technologies with a human-centric design to improve usability (C3). We validated our framework using the cable shark device assembly process, demonstrating its ability to facilitate intuitive human-robot communication via voice commands with language variations. We dynamically adapt to task variations and errors by integrating LLM, sensors, and task control mechanisms, demonstrating its ability to maintain productivity and ensure a continuous workflow.
Our next steps will test the framework under real industrial conditions, including operator variations and different manufacturing environments (e.g., noise, dust, brightness). Additionally, future work includes extending the LLM-based framework to increase adaptability by feeding the LLM with diverse data on robotic tasks and sensor information. This will enhance the robot’s task flexibility, safety, and ability to handle unexpected errors. We also plan to incorporate multiple modalities, such as haptic and gestures to improve human-robot interaction. By integrating LLM to this multimodal strategy, we aim to improve communication efficiency, task adaptability, and intuitive interaction in the manufacturing environment.